Fermi National Laboratory


Accelerator Update for April 25, 2005 - April 27, 2005

Monday April 25
The day shift began with Operation monitoring store 4111 and the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Meson MCenter experiment taking beam, and with MiniBooNE off due to a LCW problem.

Operations reported at 8:36 AM, that water technicians were beginning their repair of the MiniBooNE LCW leak.

At 10:32 AM, a Linac RF station (KRF5) tripped off. It reset okay.

At 10:41 AM, KRF6 tripped off for the first of five times in sixteen minutes. It reset okay each time.

At 2:35 PM, Operations reported that the voltage for the Linac I-Source began wandering.

Operations began sending beam to MiniBooNE at 4:29 PM.

Operations reported at 5:05 PM that Tevatron (TeV) experts began an end-of-store study.

Operation terminated store 4111 at 5:30 PM.

Operations put the TeV into shot setup at 6:12 PM.

At 6:28 PM, a Booster RF station (BRF16) tripped off. It reset okay.

Operations established combination store 4112 at 8:58 PM with an initial luminosity of 109.93E30. Pbar resumed stacking.

MCenter resumed taking beam at 9:09 PM.

NuMI experts began conducting target scans at 9:19 PM.

At 11:47 PM, the Booster West Anode power supply (WAPS) tripped off and wouldn't reset. Operators investigated and discovered that BRF16 was the cause. Operations called in an expert for help.

Tuesday April 26
The midnight shift began with Operations monitoring store 4112 and the stash, and with Booster off due to a problem with WAPS.

Operations managed to reset WAPS at 12:33 AM, but its modulator voltage was about 10% low. Booster ran at a reduced rate while waiting for BRF16 repairs. Pbar resumed stacking, and MiniBooNE, MCenter, and NuMI all resumed taking beam.

NuMI experts completed their target scans at 1:09 AM.

At 2:53 AM, BRF16 tripped off. It reset okay.

Operations put the Recycler into shot setup to stash antiprotons at 4:43 AM. Operators completed the transfers about an hour later.

MCenter resumed taking beam at 6:05 AM.

At 8:24 AM, a TeV sector F1 vacuum problem appeared. Operators investigated and found a blown fuse in a vacuum chassis. They replaced the fuse.

An RF expert brought BRF19 back on line at 9:20 AM.

Recycler began stashing antiprotons at 1:02 PM. The stash ended an hour later. The transfer efficiency was poor due to a Main Injector (MI) cogging problem. Pbar resumed stacking.

MCenter resumed taking beam at 2:16 PM.

At 4:41 PM, Operations reported that the 400 MeV chopper began missing pulses. Operators informed experts.

At 7:46 PM, LRF4 tripped off. After several failed attempts to reset the station, operators called in experts.

At 10:09 PM, BRF2 tripped off due to a LCW flow inhibit. An expert happened to be around and had the station back on line by 10:53 PM

A Linac expert repaired LRF4 and Operations reestablished Linac beam at 10:39 PM. Pbar resumed stacking, and MiniBooNE and MCenter resumed taking beam.

At 10:59 PM, BRF2 tripped off again. Experts bypassed the station.

Pbar resumed stacking at 11:32 PM.

Wednesday April 27
The midnight shift began with Operations monitoring store 4112 and the stash, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 12:05 AM, a Booster East anode power supply (EAPS) tripped off. Experts investigated.

At 12:18 AM, BRF7 & 8 tripped off due to vacuum pressure, but operators found no vacuum problem. The stations reset okay.

Booster RF technicians began working on BRF2 at 12:57 AM. They brought the station back on line at 2:20 AM.

At 2:41 AM, BRF2 tripped off. Operators managed to get the station to reset about fifteen minutes later.

Operations terminated store 4112 at 5:08 AM. TeV experts began a study period.

At 5:39 AM, BRF17 tripped off. It reset okay.

The Plan for Wednesday
The plan for today is to allow experts to conduct four hours of TeV separator conditioning. During that time both the Collider Detector at Fermilab and D-Zero experiments will access their collision halls. Cryo system techs will work on the cryo systems at TeV sectors C2 and D2. Pbar will conduct a reverse proton study. The next store will go in this afternoon.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 4/27/2005   email Fermilab